SCI Question 5

SOLVE FOR THE VALUE OF COMPREHENSIVE INCOME

 

OCI = 0

Operating Profit = 5,000,000

Income Tax Expense = 25% of Operating Profit


1
Expert's answer
2021-09-30T15:42:21-0400

"Comprehensive Income =Operating\ profit -Income \ tax \ expense\\ Income \ tax \ expense = 0.25 \times 5,000,000\\=1,250,000 \\Comprehensive Income=5,000,000-1,250,000 \\ = 3,750,000"